Consider all the possible samples of size 2 that can be drawn with replacement from the population 1, 2, 3, 4, 5, and 6. Create a sampling distribution of the sample mean


1
Expert's answer
2021-06-15T05:34:22-0400

We have population values "1,2,3,4,5,6" population size "N=6" and sample size "n=2."Thus, the number of possible samples which can be drawn with replacement is


"(N)^n=6^2=36"


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,1 & 1.0 \\
 \hdashline
 2 & 1,2 & 1.5 \\
 \hdashline
 3 & 1,3 & 2.0 \\
 \hdashline
 4 & 1,4 & 2.5\\
 \hdashline
 5 & 1,5 & 3.0 \\
\hdashline
 6 & 1,6 & 3.5 \\
 \hline
 7 & 2,1 & 1.5 \\
 \hline
 8 & 2,2 & 2.0 \\
 \hline
 9 & 2,3 & 2.5 \\
 \hline
 10 & 2,4 & 3.0 \\
 \hline
 11 & 2,5 & 3.5 \\
 \hline
12 & 2,6 & 4.0 \\
 \hline
13 & 3,1 & 2.0 \\
 \hline
14 & 3,2 & 2.5 \\
 \hline
15 & 3,3 & 3.0 \\
 \hline
16 & 3,4 & 3.5 \\
 \hline
17 & 3,5 & 4.0 \\
 \hline
18 & 3,6 & 4.5 \\
 \hline
19 & 4,1 & 2.5 \\
 \hline
20 & 4,2 & 3.0 \\
 \hline
21 & 4,3 & 3.5 \\
 \hline
 22 & 4,4 & 4.0 \\
 \hline
23 & 4,5 & 4.5 \\
 \hline
24 & 4,6 & 5.0 \\
 \hline
25 & 5,1 & 3.0 \\
 \hline
26 & 5,2 & 3.5 \\
 \hline
27 & 5,3 & 4.0 \\
 \hline
28 & 5,4 & 4.5 \\
 \hline
29 & 5,5 & 5.0 \\
 \hline
30 & 5,6 & 5.5 \\
 \hline
31 & 6,1 & 3.5 \\
 \hline
32 & 6,2 & 4.0 \\
 \hline
33 & 6,3 & 4.5 \\
 \hline
34 & 6,4 & 5.0 \\
 \hline
35 & 6,5 & 5.5 \\
 \hline
36 & 6,6 & 6.0 \\
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
1.0& 1 & 1/36 & 2/72 & 4/144 \\
 \hdashline
 1.5 & 2 & 2/36 & 6/72 & 18/144 \\
 \hdashline
 2.0 & 3 & 3/36 & 12/72 & 48/144 \\
 \hdashline
 2.5 & 4 & 4/36 & 20/72 & 100/144 \\
 \hdashline
 3.0 & 5 & 5/36& 30/72 & 180/144 \\
 \hdashline
 3.5 & 6 & 6/36 & 42/72 & 294/144 \\
 \hdashline
 4.0 & 5 & 5/36 & 40/72 & 320/144\\
 \hdashline
 4.5 & 4 & 4/36 & 36/72 & 324/144 \\
 \hdashline
 5.0 & 3 & 3/36 & 30/72 & 300/144\\
 \hdashline
 5.5 & 2 & 2/36 & 22/72 & 242/144 \\
 \hdashline
 6.0 & 1& 1/36 & 12/72 & 144/144\\
 \hdashline
 Total & 36 & 1 & 3.5 & 329/24 \\ \hline
\end{array}"



Mean


"\mu=\dfrac{1+2+3+4+5+6}{6}=3.5"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=3.5"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=3.5=\mu"



Variance


"\sigma^2=\dfrac{1}{6}\big((1-3.5)^2+(2-3.5)^2+(3-3.5)^2""(4-3.5)^2+(5-3.5)^2+(6-3.5)^2\big)=\dfrac{8.75}{3}"


Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{8.75}{3}}\approx1.7078"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{329}{24}-(\dfrac{7}{2})^2=\dfrac{35}{24}"




"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{35}{24}}\approx1.2076"


If a population is infinite and the sampling is random or if the population is finite and the sampling is with replacement, then the variance of the sampling distribution of means, "Var(\bar{X})" is given by  


"Var(\bar{X})=\dfrac{\sigma^2}{n}"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}=\dfrac{8.75}{3(2)}=\dfrac{8.75}{6}=\dfrac{35}{24}, True"